OpenStudy (anonymous):

Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). A. drawing a black seven out of a standard 52 card deck given it’s not a face card or an Ace. B. rolling a sum of 4 on two fair six sided dice C. drawing a three out of a standard 52 card deck given it’s not a face card or an Ace. D. rolling a sum of 9 on two fair six sided dice E. rolling a sum of 7 on two fair six sided dice

OpenStudy (anonymous):

what I have so far is: A=2/36 B=3/36 C=4/36 D=4/36 E=6/36

OpenStudy (anonymous):

so the 2 that have equal would be c and d

OpenStudy (anonymous):

yes. let me know if it's unclear how I got those numbers.